# Long-Context Linear System Identification

### v1

@@ B.4 Concentration inequalities
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{S}:\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant\gamma\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\leqslant\\ &\quad\quad|\mathcal{S}|\exp\left(-\dfrac{\frac{1}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}}{2C^{\prime}(\delta/4)R\gamma\ln dTN}+\ln\left(\ln\left(\dfrac{4\|L_{\star}\|_{\rm op}}{R\sigma_{\min}(L_{\star})^{2}\eta D}\right)+1\right)\right)+\frac{\delta}{2}\,.\end{split}$ (35)
Recall that Lemmas B.22, B.23 and B.21 imply
$\begin{split}\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}&\geqslant\frac{1}{2}\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}-\sigma^{2}(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}})p^{\prime}dNT\epsilon^{2}\|L_{\star}\|_{\rm op}^{2}\,,\\ \operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)&\leqslant\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E)+\sigma^{2}(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}})\sqrt{p^{\prime}}dNT\epsilon\|L_{\star}\|_{\rm op}\,,\\ \end{split}$
with probability at least $41-\delta/$ 4. We set $\epsilon$ as follows:
$\epsilon=\min{{\left\{\dfrac{\frac{1}{2}\|L_{\star}\|_{\rm op}C\eta Dp^{\prime}dr\left(\ln\tau\right)^{3/2}}{(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}})\sqrt{p^{\prime}}dNT\|L_{\star}\|_{\rm op}},\dfrac{1}{2}\sqrt{\dfrac{\frac{1}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}}{(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}})p^{\prime}dNT\|L_{\star}\|_{\rm op}^{2}}}\right\}}}\,.$
In particular, $\epsilon$ is small such that for any ${\boldsymbol{A}}_{1}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$,
$\exists{\boldsymbol{A}}_{2}\in\mathcal{S}:\quad\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}\leqslant\frac{1}{4}\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}\,,\quad\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E)\leqslant 2\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)\,.$
Then, Equation 35 implies that
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D):\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant\gamma\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\leqslant\\ &\quad\quad|\mathcal{S}|\exp\left(-\dfrac{\frac{1}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}}{16C^{\prime}(\delta/4)R\gamma\ln dTN}+\ln\left(\ln\left(\dfrac{4\|L_{\star}\|_{\rm op}}{R\sigma_{\min}(L_{\star})^{2}\eta D}\right)+1\right)\right)+\frac{3\delta}{4}\,,\end{split}$ (36)
by the following computation:
@@
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D):\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant\gamma\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\\ &\quad\quad\quad\quad\leqslant\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{S}:\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant\frac{\gamma}{8}\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\,.\\ \end{split}$
We now have to show that the right-hand side of Equation 36 for $\gamma=2$ is upper bounded by $\delta$. That is, we need to prove
@@
$\begin{split}&\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}\\ &\quad\quad\geqslant 256C^{\prime}(\delta/4)R\ln dTN\left(\ln|\mathcal{S}|+\ln\left(\ln\left(\dfrac{4\|L_{\star}\|_{\rm op}}{R\sigma_{\min}(L_{\star})^{2}\eta D}\right)+1\right)+\delta/4\right)\,.\end{split}$
In order to simplify the expressions, we set $R=4$, plug in lower bound for $\sigma_{\min}(L_{\star})$ and upper bound for $\|L_{\star}\|_{\rm op}$ from Lemma B.11, lower bound $\eta$ with $1$, upper bound $\ln dTN$ with $2\ln\tau$ and derive the following looser condition:
$C\frac{D^{2}}{(D+1)^{2}}p^{\prime}dr\ln\tau\geqslant 2048C^{\prime}(\delta/4)\left(\ln|\mathcal{S}|+\ln\left(\ln\left(\dfrac{\left(D^{T}-1\right)\left(D+1\right)^{2}}{(D-1)D}\right)+1\right)+\delta/4\right)\,.$ (37)
The cardinality of $\mathcal{S}$ is smaller than the cardinality of an $\epsilon$-net $\mathcal{S}^{\prime}$ covering all of $\mathcal{A}(D)\setminus\mathcal{G}(C,D)\subset\mathcal{A}(D)$. Therefore, we have the following upper bound on $\ln|\mathcal{S}|$:
$\ln|\mathcal{S}|\leqslant 9p^{\prime}dr\ln\frac{D}{\epsilon}\leqslant 9p^{\prime}dr\ln\left(4\sqrt{2}\left(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}}\right)NT\sigma_{{\mathrm{cond}}}(L_{\star})\right)\,,$ (38)
where we grossly upper bound $\frac{1}{\epsilon}$ similar to Equation 37. Then, Equation 37 is satisfied if:
@@
$\begin{split}Cp^{\prime}dr\ln\tau&\geqslant 2048C^{\prime}(\delta/4)\frac{(D+1)^{2}}{D^{2}}\left(18p^{\prime}dr\ln\tau+9\ln\left(4\sqrt{2}\left(1+c^{2}\dfrac{\ln\frac{4}{\delta}}{C_{HW}}\right)\right)\right)\\ &\quad+2048C^{\prime}(\delta/4)\frac{(D+1)^{2}}{D^{2}}\left(\ln\left(\ln\left(\dfrac{\left(D^{T}-1\right)\left(D+1\right)^{2}}{(D-1)D}\right)+1\right)+\delta/4\right)\,,\end{split}$ (39)
where we plug in the upper bound in Equation 38 and then bound $NT\sigma_{{\mathrm{cond}}}(L_{\star})$ with $\tau^{2}$. Now, note that there exist a constant $C^{\prime\prime}(D)$ such that
@@
$\ln\left(\ln\left(\dfrac{\left(D^{T}-1\right)\left(D+1\right)^{2}}{(D-1)D}\right)+1\right)=C^{\prime\prime}(D)\ln T\leqslant C^{\prime\prime}(D)\ln\tau\,.$
Therefore, Equation 39 is satisfied for a constant $C=\mathcal{O}(\ln\frac{1}{\delta})$.
Finally, observe that the dependencies of $C$ on $c$ are due to the noise concentration inequalities and are the results of applications of Lemma B.21. Therefore, one can completely remove this dependency by fixing $\delta>\exp(-\nu dNT)+\delta^{\prime}$ where $\nu=\nu(c)$ is a well-chosen constant and $\delta^{\prime}>0$ is small and arbitrary. The $\ln\ln D$ dependency, however, is due to the martingale concentration inequalities in Theorem B.26. Therefore, the dependency of $C$ on $D$ is inherent to the process and cannot be removed.
Appendix C Experiments
@@
For all the experiments, ${\boldsymbol{A}}^{\star}$ is generated as follows. First, $p$ orthogonal matrices of shape $d\times d$ are sampled. These are then scaled down by $\alpha\cdot p$ where $\alpha$ is arbitrarily set to $0.5$. In cases where ${\boldsymbol{A}}$ needs to be initialized, we use the same recipe for the student model with $p^{\prime}$ instead of $p$ and set $\alpha=1$. For experiments with low-rank ground truth, we set arbitrary $d-r$ singular values to 00 following a SVD decomposition. Each experiment in this section has been run over 3 independent seeds and the average is plotted. As the variance is small and the plots usually overlap, we opt to not plot it for visual clarity.
Theorems 4.1, 4.2 and 4.3 provide rates on estimation error for empirical minimizers. In the following, we study these rates empirically for various values of $p^{\prime},p,d,N,T$ and $r$ where $r=d$ for full-rank experiments or $r=5$ for low-rank experiments and $p^{\prime}=p$ except it is stated otherwise. We use two quantities, $\beta=NT$, the number of total tokens, $\gamma=pdr$, the number of parameters to estimate, to summarize information in the plots. For Theorems 4.1 and 4.3, $\hat{A}$ is computed with the OLS estimator and for Theorem 4.2, $\hat{A}$ is learned with gradient descent with learning rate $\alpha$ on the group-norm regularized loss in Equation 10. The parameter $\lambda$ and learning rate $\alpha$ are tuned by a grid search.
@@
Figure 1: Scaling of estimation error with respect to the ratio $\beta/\gamma=NT/pd^{2}$ with the OLS estimator. The black dashed line plots $\gamma/\beta$ for reference.
Figure 2: Scaling of estimation error for different values of $p$, $d$ and $N$ with the OLS estimator. Recall that $\beta=NT,\gamma=pd^{2}$ and $\bar{\beta}=\beta/\ln(1+\sqrt{N})$. Black dashed lines are drawn for reference and equals to $\sqrt{\gamma/\beta}$.
Figure 3 plots the estimation error for different degrees of misspecification where the context length is fixed to $p=15$. The curves for various $p^{\prime}\in{{\left\{5,10,15\right\}}}$ overlap, which verify the rate $\gamma/\beta=p^{\prime}d^{2}/NT$ predicted by Theorem 4.3 holds.
Figure 3: Scaling of estimation error with respect to the ratio $\beta/\gamma=NT/p^{\prime}d^{2}$ for different $p^{\prime}=5,10,15$ with the OLS estimator. The black dashed line plots $\gamma/\beta$ for reference.
Figure 4 repeats the same plots for low-rank experiments where $d=15,r=5$ are fixed and $p,N$ and $T$ are varied as before. Good estimation of ${\boldsymbol{A}}$ is not straightforward as $\lambda$ has to be appropriately tuned. Yet, we see that the group-nuclear norm regularized estimators found with gradient descent after tuning on regularization problem $\lambda\in{{\left\{10^{-1},10^{-2},10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\right\}}}$ and learning rate $\alpha\in{{\left\{10^{-1},10^{-2},10^{-3}\right\}}}$ obtain improved estimation errors than non-regularized OLS estimator. Particularly, the sample efficiency benefits of the group-nuclear norm regularization are amplified in the low-data regime. We leave the analysis of group-nuclear norm regularization as a future work.

### v2

@@ C.4 Concentration inequalities
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D):W_{{\boldsymbol{A}}}\leqslant 2\sigma^{2}\eta Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}\right)&\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant 16\sigma^{2}\eta Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}\right)\,,\end{split}$
where $\mathcal{E}$ is the event that both Equation 26 and Equation 27 hold.
By Theorem C.4, there exists a universal constant $C^{\prime}$ such that
@@
($\begin{split}&\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{\boldsymbol{A}}\leqslant 16\sigma^{2}\eta Y_{\boldsymbol{A}}\right\}}}\cap\mathcal{E}\right)\leqslant\\ &\quad\quad\quad\quad|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}+\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)+\frac{\delta}{4}\,,\end{split}$ (29)
with the following choices:
@@
$\begin{split}\gamma&=16\sigma^{2}\eta\,,\quad R=\gamma^{\prime}=C^{\prime}e\gamma\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)\,,\\ \alpha_{L}&=\frac{\sigma^{4}}{8}\sigma_{\min}(L_{\star})^{2}CD^{2}\eta^{2}\tau p^{\prime}dr\,,\\ \alpha_{U}&=\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{CD^{2}\eta^{2}\tau p^{\prime}d^{2}rNT}\,.\end{split}$
By a union bound, Equation 29 implies that
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D):W_{{\boldsymbol{A}}}\leqslant 2\sigma^{2}\eta Y_{{\boldsymbol{A}}}\right)\leqslant\\ &\quad\quad\quad\quad|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}+\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)+\frac{3\delta}{4}\,.\end{split}$ (30)
By Lemma C.4, we have
$\|\Delta_{{\boldsymbol{A}}}\|_{\rm op}=\|M_{{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}}\|_{\rm op}\geqslant\|{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\|_{\rm op}\,.$
Therefore, an $\epsilon$-net covering of $\mathcal{M}_{r,p^{\prime}}(D)$ defined in Corollary B.2 is also an $\epsilon$-net over $\mathcal{A}(D)$ after a shift by ${\boldsymbol{A}}^{\star}$ and we have
$\ln|\mathcal{S}|\leqslant 9p^{\prime}dr\ln\dfrac{Dp^{\prime}}{\epsilon}\,.$
Finally, we have to show that the right-hand side of Equation 30 is upper bounded by $\delta$. That is, we need to prove
@@
$\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}\geqslant\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)+9p^{\prime}dr\ln\dfrac{Dp^{\prime}}{\epsilon}+\ln\dfrac{4}{\delta}\,,$ (31)
for a suitable choice of $C$. Note that Equation 31 is homogeneous in $\sigma$. In addition, the left-hand side does not depend on $\eta$, whereas the right-hand side is decreasing in $\eta$. Therefore, for simplicity, we can set $\sigma=1$ and $\eta=1$.
By Lemma C.3 we have
$\sigma_{\min}(L_{\star})\geqslant\dfrac{1}{D+1}\geqslant\dfrac{1}{2D}\,,$
and
$\alpha_{L}\geqslant\dfrac{C\tau p^{\prime}dr}{32}=\Omega\left(C\left(\left(1+\ln p^{\prime}dNT\right)^{3}\left(1+\ln\kappa\right)\right)p^{\prime}dr\right)\,.$
For a choice that verifies
$C=\Theta\left(\left(1+\ln\dfrac{1}{\delta}\right)^{3}\right)\,,$ (32)
we have the following lower bound for $\alpha_{L}$:
$\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}=\Omega\left(\left(1+\ln p^{\prime}dNT\right)\left(1+\ln\kappa\right)\left(1+\ln\dfrac{1}{\delta}\right)p^{\prime}dr\right)\,.$ (33)
The condition Equation 31 is satisfied if the following three conditions are individually satisfied:
(i) $\quad\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}\geqslant\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)\,,$
(ii) $\quad\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}\geqslant 9p^{\prime}dr\ln\dfrac{Dp^{\prime}}{\epsilon}\,,$
(iii) $\quad\dfrac{\alpha_{L}}{4\gamma^{\prime 2}}\geqslant\ln\dfrac{4}{\delta}\,,$
as the left-hand side of Equation 31 is increasing in $C$, while the right-hand side is decreasing in $C$, we can pick the maximum $C$ that satisfies all three conditions.
Condition (i).
By Lemma C.3, we have
$2D\|L_{\star}\|_{\rm op}\geqslant(D+1)\|L_{\star}\|_{\rm op}\geqslant(D+1)\sigma_{\min}(L_{\star})\geqslant 1\,,$
and
$\dfrac{\alpha_{U}}{\alpha_{L}}\leqslant\dfrac{2D\|L_{\star}\|_{\rm op}\alpha_{U}}{\alpha_{L}}=\mathcal{O}\left(\sqrt{NT}\kappa^{2}\right)\,.$
As $x\geqslant\ln x+1$ for any $x\geqslant 0$, we have
$\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)\leqslant\ln\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}=\mathcal{O}\left(\ln 2dNT+\ln\kappa+\ln\dfrac{1}{\delta}\right)\,.$
The lower bound in Equation 33 is sufficient to satisfy this condition.
Condition (ii).
We have that
$\begin{split}\ln\dfrac{Dp^{\prime}}{\epsilon}&\leqslant\ln Dp^{\prime}\max{{\left\{\sqrt{\dfrac{\epsilon_{1}(\delta/4)}{\alpha_{L}}},\dfrac{4\epsilon_{2}(\delta/4)}{\alpha_{L}},\dfrac{1}{D}\right\}}}\\ &\leqslant\ln Dp^{\prime}\max{{\left\{\dfrac{D\epsilon_{1}(\delta/4)}{\alpha_{L}},\dfrac{4\epsilon_{2}(\delta/4)}{\alpha_{L}},\dfrac{1}{D}\right\}}}\\ &=\mathcal{O}\left(\ln p^{\prime}NT+\ln\kappa+\ln\dfrac{1}{\delta}\right)\,.\end{split}$
The lower bound in Equation 33 is sufficient to satisfy this condition.
Condition (iii).
This condition is readily verified by Equation 33.
Remark C.3.
The choice of $C$ in Equation 32 and $\tau$, which appear in the bounds, have third-order logarithmic dependencies on $\delta$ and the problem parameters, respectively. These arise from the combination of a high-probability bound on the sub-Gaussian noise in terms of $\delta$ and the problem parameters, and a union bound over a discretization net. Assuming strictly bounded noise, i.e., $\sup\|\xi_{t}^{(n)}\|_{\infty}\leqslant B$ for some $B>0$, would yield an improved rate with respect to both $\delta$ and problem parameters, and a tighter analysis by employing chaining techniques could potentially reduce the logarithmic dependencies in problem parameters.
Appendix D Approximate empirical risk minimization
In this section, we extend the results of Appendix C to approximate empirical risk minimizers ${\boldsymbol{\tilde{A}}}$ as in Equation 12. We begin with trivial extensions of Proposition C.2 and Corollary C.4 to approximate minimizers.
Proposition D.1.
Let ${\boldsymbol{\tilde{A}}}$ be an estimate that provides an $\epsilon_{\mathrm{tr}}$-approximation to the loss of $\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})$:
$\mathcal{L}({\boldsymbol{\tilde{A}}})\leqslant\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})+\epsilon_{\mathrm{tr}}\,.$
In particular, one can choose ${\boldsymbol{\tilde{A}}}$ as an $\epsilon_{\mathrm{tr}}$-approximate empirical risk minimizer, i.e.,
${\boldsymbol{\tilde{A}}}\in{{\left\{{\boldsymbol{A}}\in\mathcal{A}(\mathcal{D})\mid\mathcal{L}({\boldsymbol{A}})\leqslant\mathcal{L}({\boldsymbol{\hat{A}}})+\epsilon_{\mathrm{tr}}\right\}}}\,,$
where ${\boldsymbol{\hat{A}}}$ is the empirical risk minimizer defined in Equation 21. Then, ${\boldsymbol{\tilde{A}}}$ satisfies
$\begin{split}{{\left\|\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}^{\top}\left(I-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right)+NT\epsilon_{\mathrm{tr}}\,.\end{split}$
Proof.
The proof follows from Equation 22. ∎
Corollary D.1.
Observe that for $p^{\prime}=p$, Proposition D.1 reads
${{\left\|\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right)+NT\epsilon_{\mathrm{tr}}\,.$
For $p^{\prime}<p$, one can write the following relaxed condition for any ${\boldsymbol{\tilde{A}}}$:
$\begin{split}{{\left\|\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right\|}}_{F}^{2}&\leqslant 2\eta\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right)+NT\epsilon_{\mathrm{tr}}\,.\end{split}$
Proof.
The proof follows from Corollary C.4. ∎
We divide the analysis into two cases: $2\eta\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right)>NT\epsilon_{\mathrm{tr}}$ and $2\eta\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right)\leqslant NT\epsilon_{\mathrm{tr}}$. In the first case, we can directly apply the results of Appendix C to approximate minimizers after noting that Corollary C.4 implies that
${{\left\|\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\left(2\eta\right)\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right)\,.$
This is equivalent to doubling the constant $\eta$ in the main theorem and does not change our results.
In the second case, we have the following:
${{\left\|\Delta_{{\boldsymbol{\tilde{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2NT\epsilon_{\mathrm{tr}}\,.$
Recall the lower isometry proven in Equation 26:
$\inf_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}CD^{2}\eta^{2}\tau p^{\prime}dr\,.\$
If $\epsilon_{\mathrm{tr}}$ is such that
$2NT\epsilon_{\mathrm{tr}}\leqslant\frac{\sigma^{4}}{8}\sigma_{\min}(L_{\star})^{2}CD^{2}\eta^{2}\tau p^{\prime}dr\,,$
then the approximate minimizer ${\boldsymbol{\tilde{A}}}$ is guaranteed to be in the set $\mathcal{G}(C,D)$ with high probability. By Corollary C.3, $\mathcal{G}(C,D)\subset\mathcal{S}(C,D)$, and thus, we have the desired result ${\boldsymbol{\tilde{A}}}\in\mathcal{S}(C,D)$.
Appendix E Experiments
@@
For all the experiments, ${\boldsymbol{A}}^{\star}$ is generated as follows. First, $p$ orthogonal matrices of shape $d\times d$ are sampled. These are then scaled down by $\alpha\cdot p$ where $\alpha$ is arbitrarily set to $0.5$. In cases where ${\boldsymbol{A}}$ needs to be initialized, we use the same recipe for the student model with $p^{\prime}$ instead of $p$ and set $\alpha=1$. For experiments with low-rank ground truth, we set arbitrary $d-r$ singular values to 00 following a SVD decomposition. Each experiment in this section has been run over 3 independent seeds and the average is plotted. As the variance is small and the plots would otherwise overlap, we have opted not to plot it for visual clarity.
Theorems 4.1, 4.2 and 4.3 provide rates on estimation error for empirical minimizers. In the following, we study these rates empirically for various values of $p^{\prime},p,d,N,T$ and $r$ where $r=d$ and $p^{\prime}=p$ unless stated otherwise. We use two quantities, $\beta=NT$, the number of total tokens, $\gamma=pdr$, the number of parameters to estimate, to summarize information in the plots. For Theorems 4.1 and 4.3, $\hat{A}$ is computed with the OLS estimator and for Theorem 4.2, $\hat{A}$ is learned with gradient descent with learning rate $\alpha$ on the group-norm regularized loss in Equation 9. The parameter $\lambda$ and learning rate $\alpha$ are tuned by a grid search.
@@
Figure 1: Scaling of estimation error with respect to the ratio $\beta/\gamma=NT/pd^{2}$ with the OLS estimator. The black dashed line marks the reference value $\gamma/\beta$.
Figure 2: Scaling of estimation error for different values of $p$, $d$ and $N$ with the OLS estimator. Recall that $\beta=NT,\gamma=pd^{2}$ and $\bar{\beta}=\beta/\ln(1+\sqrt{N})$. The black dashed lines mark the reference values corresponding to $\gamma/\beta$ and $\gamma/\bar{\beta}$.
Figure 3 plots the estimation error for different degrees of misspecification where the context length is fixed to $p=15$. The curves for various $p^{\prime}\in{{\left\{5,10,15\right\}}}$ overlap, which verify that the rate $\gamma/\beta=p^{\prime}d^{2}/NT$ predicted by Theorem 4.3 holds.
Figure 3: Scaling of estimation error with respect to the ratio $\beta/\gamma=NT/p^{\prime}d^{2}$ for different $p^{\prime}=5,10,15$ with the OLS estimator. The black dashed line marks the reference value $\gamma/\beta$.
Figure 4 repeats the same plots for low-rank experiments where $d=15,r=5$ are fixed and $p,N$ and $T$ are varied as before. Good estimation of ${\boldsymbol{A}}$ is not straightforward as $\lambda$ has to be appropriately tuned. Yet, we see that the group-nuclear norm regularized estimators found with gradient descent after tuning on regularization problem $\lambda\in{{\left\{10^{-1},10^{-2},10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\right\}}}$ and learning rate $\alpha\in{{\left\{10^{-1},10^{-2},10^{-3}\right\}}}$ obtain lower estimation errors than non-regularized OLS estimator. Particularly, the sample efficiency benefits of the regularization are amplified in the low-data regime. We leave the analysis of group-nuclear norm regularization as a future work.